OpenStudy (anonymous):

[8.07] Which polynomial is a difference of two squares? x2 − 16 x2 + 16 x2 − 8 x2 + 8

jimthompson5910 (jim_thompson5910):

difference of squares is of the form a^2 - b^2

OpenStudy (anonymous):

what is next

jimthompson5910 (jim_thompson5910):

is x^2 + 8 a difference of squares?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

why not?

OpenStudy (anonymous):

x^2 − 25 = x^2 - 5^2 = (x-5)(x+5)

jimthompson5910 (jim_thompson5910):

that's a difference of squares, yes, but that's not an answer choice

jimthompson5910 (jim_thompson5910):

hopefully you got the idea down

OpenStudy (anonymous):

will it be B

jimthompson5910 (jim_thompson5910):

that's a sum of squares

jimthompson5910 (jim_thompson5910):

you want a difference of squares

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

How about C

jimthompson5910 (jim_thompson5910):

is 8 a perfect square?

OpenStudy (anonymous):

yea

jimthompson5910 (jim_thompson5910):

how so?

jimthompson5910 (jim_thompson5910):

what whole number squared gives you 8?

jimthompson5910 (jim_thompson5910):

25 is a perfect square since 5^2 = 25

OpenStudy (anonymous):

its no